OpenStudy (anonymous):

12 miles/27 hours = 4 miles/h hours 28 12 9 18

OpenStudy (alexandervonhumboldt2):

\[\frac{ 12 }{ 27 }=\frac{ 4 }{ h }\]

OpenStudy (mrnood):

this means \[\frac{ 12m }{ 27h } = \frac{ 4m }{ xh }\]

OpenStudy (alexandervonhumboldt2):

cross multiply: \[12h=4*27\]

OpenStudy (alexandervonhumboldt2):

now solve for h.

OpenStudy (anonymous):

i know i probably should but i still dont get it im just dumb:(

OpenStudy (kobeni-chan):

If you look at the steps @AlexandervonHumboldt2 gave, you will take the value of 4 times 27, and then divide that number by 12 (from 12h), so you can get h by itself.

OpenStudy (anonymous):

oooh thanks i didnt understand it at first

OpenStudy (kobeni-chan):

That's ok! :)

OpenStudy (anonymous):

so if im right 9

OpenStudy (kobeni-chan):

Yes :)

OpenStudy (anonymous):

Yes! thank you.

OpenStudy (kobeni-chan):

No problem!